OpenStudy (anonymous):

use the distributive property 7(w-2)=-14 is the answer -2 if dividing by 7

OpenStudy (anonymous):

w=0

OpenStudy (anonymous):

are you sure

OpenStudy (anonymous):

7w-14=-14 7w=0 w=0

OpenStudy (anonymous):

divide by 7 and get w-2=-2 then add 2 and get w=0

OpenStudy (anonymous):

using distributive law it becomes 7w - 14 = -14 7w = -14 + 14 7w = 0 w = 0/7w = 0

OpenStudy (anonymous):

thnks all of you

OpenStudy (anonymous):

welcome...